# Test-time regression: a unifying framework for designing sequence models with associative memory

### v1

@@ Abstract
Sequences provide a remarkably general way to represent and process information. This powerful abstraction has placed sequence modeling at the center of modern deep learning applications, inspiring numerous architectures from transformers to recurrent networks. While this fragmented development has yielded powerful models, it has left us without a unified framework to understand their fundamental similarities and explain their effectiveness. We present a unifying framework motivated by an empirical observation: effective sequence models must be able to perform associative recall. Our key insight is that memorizing input tokens through an associative memory is equivalent to performing regression at test-time. This regression-memory correspondence provides a framework for deriving sequence models that can perform associative recall, offering a systematic lens to understand seemingly ad-hoc architectural choices. We show numerous recent architectures — including linear attention models, their gated variants, state-space models, online learners, and softmax attention — emerge naturally as specific approaches to test-time regression. Each architecture corresponds to three design choices: the relative importance of each association, the regressor function class, and the optimization algorithm. This connection leads to new understanding: we provide theoretical justification for QKNorm in softmax attention, and we motivate higher-order generalizations of softmax attention. Beyond unification, our work unlocks decades of rich statistical tools that can guide future development of more powerful yet principled sequence models.
@@
This development has produced a diversity of architectures, each with its own unique characteristics and performance trade-offs. While these architectures have achieved considerable success, they have largely emerged through separate lines of investigation. This fragmented and often empirically-driven approach to model development limits our ability to systematically understand and improve design choices. Moreover, the idiosyncratic notations of each architecture obscures their underlying connections (Rush, 2024). Given the wide variety of sequence models, from transformers to convolutional networks to recurrent networks, a natural question to ask is whether there is an underlying principle that explains why some sequence models work better than others.
One empirical discovery that ties together disparate architectures is the strong correlation between their associative recall ability and their language modeling performance (Olsson et al., 2022, Arora et al., 2023). For example, Olsson et al. (2022) found that transformers learn induction heads that store consecutive tokens as key-value pairs to predict future tokens. Given the empirical importance of associative recall, how can we systematically design architectures that can perform associative recall? In this paper, we introduce a simple but principled framework that results in sequence models with associative memory, unifying broad classes of models.
Our crucial observation is that we can memorize key-values pairs by solving a regression problem (Kohonen, 1972). A sequence layer that memorizes earlier tokens for later retrieval is one that performs test-time regression over the input tokens; we name this “test-time regression”, a reference to the test-time training paradigm (Sun et al., 2024, 2020). By establishing the equivalence between associative memory and regression, we create a framework under which we can holistically understand ad-hoc architecture developments and motivate new advances. In particular, our framework allows us to repurpose well-understood regression tools, including weighted least-squares regression, online algorithms for regression, and nonparametric regression to explain existing architectures. Under our framework, a sequence layer is a mathematical consequence of choosing a regression objective, a regressor parameterization, and an optimization algorithm.
To illustrate the generality of our framework, we concisely derive linear attention (Katharopoulos et al., 2020), its feature-mapped variants (e.g. (Peng et al., 2020, Qin et al., 2021, Kasai et al., 2021, Zhang et al., 2023, Aksenov et al., 2024, Chen et al., 2024)), its gated variants (Sun et al., 2023, Orvieto et al., 2023, Katsch, 2024, De et al., 2024, Qin et al., 2024, Peng et al., 2024, Yang et al., 2024b, Beck et al., 2024), state-space models (Gu and Dao, 2024, Dao and Gu, 2024), fast-weights programmers (Schlag et al., 2021, Yang et al., 2024a, a), online learners (Liu et al., 2024, Sun et al., 2024, Yang et al., 2024a, Behrouz et al., 2024), and even softmax self-attention (Vaswani et al., 2017) as points within our design space. Past work by Schmidhuber (1992), Schlag et al. (2021), Clark et al. (2022), Liu et al. (2024), Yang et al. (2024a), Sun et al. (2024) proposed test-time training/fast-weight architectures that perform gradient descent at test time. Concurrent work by Behrouz et al. (2024) also points out the relationship between online learners and memorization. In contrast, we show that test-time regression is an even more fundamental principle that can also derive other classes of sequence layers, including the aforementioned gated linear attention models, state space models, softmax attention, and linear regression layers (von Oswald et al., 2023, Garnelo and Czarnecki, 2023). Our framework abstracts away practically important but low-level details in exchange for a clearer perspective on what these architectures have in common, enabling direct comparison. The regression-memory correspondence unlocks decades of advances in classical regression theory (Hastie et al., 2009, Green and Silverman, 1993, Wasserman, 2006), offering a systematic path toward sequence architectures that are both more powerful and theoretically sound.
Figure 1: Our test time regression framework unifies existing sequence models through the regression-memory correspondence. The equivalence between in-context regression and associative memory leads to a systematic approach for sequence model design through three key choices: the relative importance of each association, the regressor function class, and the optimization algorithm. The framework encompasses models ranging from classical linear regression with exact solvers (left) to first-order optimization approaches (middle), to nonlinear regressors (right). This unification reveals that seemingly different architectural innovations (shown in italics) are instances of well-studied concepts from regression theory but implemented as a neural network layer.
Section 1.1 Outline of our paper
In Section 2, we introduce our test-time regression framework and formalize the connection between associative memory and regression. We show how this perspective provides a systematic approach to sequence model design through three key choices: regression objective, function class, and optimization algorithm. Section 3 demonstrates the broad applicability of our framework by deriving state-of-the-art sequence architectures from first principles. In the process, we explain the effectiveness of QKNorm in softmax attention, and we derive higher-order generalizations of softmax attention, among other contributions. Section 4 examines how to construct effective key-value pairs for associative recall. Finally, Section 5 discusses the broader implications of viewing sequence models through the lens of regression, including future directions for new architectures.
@@
Section 2 Test-time regression as a framework for model design
The goal of sequence modeling is to transform a sequence of input tokens $\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$ into output tokens $\mathbf{y}_{1},\ldots,\mathbf{y}_{T}$ (Sutskever et al., 2014). Although many sequence layers exist, past works have shown that associative recall ability is crucial to this task by enabling in-context learning (Olsson et al., 2022). Though many sequence layers exist, it is not a priori clear which ones are able to perform associative recall without empirical evaluations (Arora et al., 2023). A common design choice is to transform the input tokens into key-value pairs $(\mathbf{k}_{1},\mathbf{v}_{1}),\ldots,(\mathbf{k}_{T},\mathbf{v}_{T})$ and queries $\mathbf{q}_{1},\ldots,\mathbf{q}_{T}$ before mixing them, following the success of transformers (Vaswani et al., 2017). While transforming inputs into query-key-value tokens has proven empirically effective, the field has lacked theoretical principles for designing sequence layers with associative memory capabilities. We address this gap through our test-time regression framework, which provides a principled recipe for deriving sequence architectures with mathematically grounded recall abilities.
Associative memory as regression.
@@
From our everyday experience, we already have an intuition for what an associative memory should behave like; for example, hearing a friend’s name should trigger a mental impression of that friend. This pairing is called a “cue” and a “response”. To reflect terminology in deep learning (Vaswani et al., 2017), we will refer to cues as “keys” and responses as “values”. Given a set of associations $(\mathbf{k}_{1},\mathbf{v}_{1}),\ldots,(\mathbf{k}_{T},\mathbf{v}_{T})$, an associative memory system is a system that performs associative recall: return $\mathbf{v}_{i}$ when given $\mathbf{k}_{i}$. Such a mapping between different sets of related objects is known as a hetero-associative memory in classic signal processing and neurocomputing literature (Hinton and Anderson, 1989).
With this intuition, we can now define a mathematical model for associative memory. Given key-value pairs $(\mathbf{k}_{1},\mathbf{v}_{1}),\ldots,(\mathbf{k}_{T},\mathbf{v}_{T})\in\mathbb{R}^{D_{k}}\times\mathbb{R}^{D_{v}}$ to be memorized, an associative memory system is a function ${m}:\mathbb{R}^{D_{k}}\to\mathbb{R}^{D_{v}}$ such that ${m}(\mathbf{k}_{t})\approx\mathbf{v}_{i}$ for $t=1,\ldots,T$ (Kohonen, 1972, 1989). When ${m}$ is flexible enough to interpolate all values such that ${m}(\mathbf{k}_{t})=\mathbf{v}_{t}$ for all $t$, it is a perfect associative memory system. This mapping implements a form of fuzzy memory: querying with $\mathbf{q}$ that approximates or is a corrupted version of some key $\mathbf{k}_{i}$ will retrieve a value close to $\mathbf{v}_{i}$, making the memory robust to input noise and perturbations (Hinton and Anderson, 1989).
@@
Our key observation is that finding a memory map that satisfies these constraints reduces to finding a vector-valued regressor ${m}$. We can derive an associative memory system by solving the weighted regression problem
$\operatorname*{minimize}\limits_{{m}\in\mathcal{M}}\ \ \frac{1}{2}\sum_{i=1}^{T}\gamma_{i}\lVert\mathbf{v}_{i}-{m}(\mathbf{k}_{i})\rVert_{2}^{2},$ (1)
@@
where the importance of each association is controlled by adjusting the relative weights $\{\gamma_{i}\}_{i=1}^{T}$ Classic works on associative memory are restricted to linear memory maps, corresponding to linear regression (Kohonen, 1989, Chapter 6.6), but here we allow $\mathcal{M}$ to be any function class. We refer to this relationship between regression and associative memory as the “regression-memory correspondence”.
Having established regression as a way to implement associative memory, we design sequence layers by solving Equation 1 to construct memory maps ${m}_{A}$ that satisfy ${m}_{A}(\mathbf{k}_{i})\approx\mathbf{v}_{i}$ for associations $A=\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i\in I}$. For non-causal tasks (e.g. vision transformers), we construct a single memory ${m}_{A}$ over all key-value pairs and compute outputs as $\mathbf{y}_{t}={m}_{A}(\mathbf{q}_{t})$. For causal tasks (e.g. language models), we construct separate memories ${m}_{t}$ for each prefix $A_{t}=\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i=1}^{t}$ and compute outputs as $\mathbf{y}_{t}={m}_{t}(\mathbf{q}_{t})$. The forward pass thus solves Equation 1 to obtain ${m}$ before applying it to queries. This naturally extends to multihead architectures (Vaswani et al., 2017, Shazeer, 2019, Ainslie et al., 2023), which correspond to parallel associative memories that may share keys, values, or queries.
We call such layers “test-time regression layers”, referencing the similar idea of test-time-training (Sun et al., 2020) but specialized to associative memory. Though solving Equation 1 can be seen as a self-supervised test-time training objective that regresses one view upon another (Sun et al., 2024), our associative memory perspective is more intuitive given its importance in in-context learning (Arora et al., 2023, Olsson et al., 2022).
A recipe for designing your own architecture.
@@
Although Equation 1 succinctly summarizes our framework, it leaves many details unspecified. For example, over which class of regressors $\mathcal{M}$ should we consider? Additionally, which associations should we prioritize memorizing? Finally, which optimization procedure should we use to minimize the loss? The regression-memory correspondence thus allows for a vast design space. We partition this wide landscape into three simple design choices:
@@
In this section, we consider causal sequence models that construct a new ${m}_{t}$ at each timestep using associations $A_{t}=\{(\mathbf{k}_{1},\mathbf{v}_{1}),\ldots,(\mathbf{k}_{t},\mathbf{v}_{t})\}$. The case for non-causal sequence models easily generalizes, following our previous discussion.
To demonstrate how our framework unifies our understanding of sequence models, we succinctly derive a swath of architectures commonly studied today. We categorize existing sequence layers into four distinct classes: linear attention (possibly with feature maps), linear attention with gating (which includes state-space models like Mamba), online learners, and softmax self-attention.

### v3

@@ Abstract
Sequence models lie at the heart of modern deep learning. However, rapid advancements have produced a diversity of seemingly unrelated architectures, such as Transformers and recurrent alternatives. In this paper, we introduce a unifying framework to understand and derive these sequence models, inspired by the empirical importance of associative recall, the capability to retrieve contextually relevant tokens. We formalize associative recall as a two-step process, memorization and retrieval, casting memorization as a regression problem. Layers that combine these two steps perform associative recall via “test-time regression” over its input tokens. Prominent layers, including linear attention, state-space models, fast-weight programmers, online learners, and softmax attention, arise as special cases defined by three design choices: the regression weights, the regressor function class, and the test-time optimization algorithm. Our approach clarifies how linear attention fails to capture inter-token correlations and offers a mathematical justification for the empirical effectiveness of query-key normalization in softmax attention. Further, it illuminates unexplored regions within the design space, which we use to derive novel higher-order generalizations of softmax attention. Beyond unification, our work bridges sequence modeling with classic regression methods, a field with extensive literature, paving the way for developing more powerful and theoretically principled architectures.
@@
This development has produced a diversity of architectures, each with its own unique characteristics and performance trade-offs. While these architectures have achieved considerable success, they have largely emerged through separate lines of investigation. Such a fragmented and often empirically-driven approach to model development limits our ability to systematically understand and improve design choices. Moreover, the idiosyncratic notations of each architecture obscures their underlying connections (Rush, 2024). Given the wide variety of sequence models, a natural question to ask is whether there is an underlying principle that explains why some sequence models work better than others.
One empirical discovery that ties together disparate architectures is the strong correlation between an architecture’s associative recall ability and its language modeling performance (Olsson et al., 2022, Arora et al., 2023a). Associative recall is the act of retrieving contextually relevant information based on an association with a query. (Arora et al., 2023a) gives the following example of how associative memory and retrieval can improve language modeling. Consider the example sentence from Arora et al. (2023b): “Hakuna Matata, it means no worries for the rest of your days. Hakuna           ”. To predict the next word, we can first memorize the previous occurrence of “Hakuna” and its associated value “Matata”; the next time we encounter “Hakuna” again we can retrieve its associated value from memory as our prediction. Notice that we can perform this task using only in-context information and we can make an accurate prediction even if we have never encountered this strange phrase before. Indeed, transformer-based language models have been discovered to exhibit this kind of behavior via “induction heads” that emerge during training (Olsson et al., 2022).
Given the empirical importance of associative recall, how can we systematically design neural network layers that can perform associative recall (AR)? In this paper, we introduce a simple but principled framework for deriving sequence layers designed to perform associative recall, which we call “test-time regression layers”. From the “Hakuna Matata” example, we see that associative recall has two-steps: memorization and retrieval. Our crucial observation is that we can implement the memorization step by solving a weighted regression problem. We can then generate an output by applying the regressor to a cue/query token, retrieving the most relevant token from our associative memory. Combining both memorization and retrieval into the forward pass of a sequence layer results in a layer that performs regression over the input tokens, a procedure that we call “test-time regression”, visualized in Figure 1. Our terminology reflects that the regressor is regenerated with each forward pass and depends only on the input tokens rather than a fixed training dataset.
Since any regression method can be used, our framework provides a general recipe to derive a large class of sequence layers. Under our framework, an AR-based sequence layer is a mathematical consequence of choosing (1) the regression weights, (2) the regressor parameterization, and (3) the optimization algorithm for finding the regressor. In this paper, we consider a few choices of these three “ingredients”, and show how to derive many recently proposed classes of sequence layers, illustrating the generality of our framework. Our derivations reveal that linear attention (Katharopoulos et al., 2020), its feature-mapped variants (e.g. (Peng et al., 2020, Qin et al., 2021, Kasai et al., 2021, Zhang et al., 2023a, Aksenov et al., 2024, Chen et al., 2024)), its gated variants (Sun et al., 2023, Orvieto et al., 2023, Katsch, 2024, De et al., 2024, Qin et al., 2024, Peng et al., 2024, Yang et al., 2024b, Beck et al., 2024), state-space model layers (Gu and Dao, 2024, Dao and Gu, 2024), fast-weights layers (Schlag et al., 2021, Yang et al., 2024a, a), online learning layers (Liu et al., 2024, Sun et al., 2024, Yang et al., 2024a, Behrouz et al., 2024), and softmax self-attention (Vaswani et al., 2017) are all test-time regression layers, implicitly performing memorization followed by retrieval in their forward passes, despite being developed from different perspectives. Figure 1 previews how existing architectures are instantiations of test-time regression layers within our framework.
Our derivations also lead to new understandings of existing sequence layers. We show that layers based on linear attention underperform because they fail to account for the correlation between tokens. We also show that query-key normalization, an important technique in stabilizing the training of large language models (Dehghani et al., 2023, Wortsman et al., 2023), is mathematically necessary to ensure that softmax self-attention is a proper local constant regressor. Finally, we propose a higher order generalization of softmax attention, motivated by local linear regression.
Outline of our paper.
We start in Section 2 by examining a few of the most prominent classes of sequence layers. Although each class of layers was developed from distinct motivations, their similar computation pattern hints at a common unifying theme that underlies their effectiveness. We then introduce our test-time regression framework in Section 3 to formalize the connection between associative memorization and regression. Our perspective provides a systematic approach to designing sequence layers that produce its outputs via associative recall. In Section 4, we show that indeed all of the aforementioned classes of sequence layers can be understood from a single unified perspective using the principles of test-time regression and associative recall. We demonstrate the generality of our framework by deriving these layers simply by varying how we minimize the regression objective. Section 5 empirically validates that test-time regression layers implicitly perform regression over its input tokens with a single forward pass. Then, in Section 6 we examine how to construct effective key-value pairs for associative recall in next-token prediction tasks. We discuss prior work related to associative recall and memory in Section 7. We finish in Section 8 by discussing the broader implications of viewing sequence models through the lens of regression, including future directions for new architectures.
Figure 1: Our framework provides a systematic way to derive sequence models that can perform associative recall, following a two step process of memorization and retrieval. The memorization step can be formalized as a solving a regression problem at test-time. Our perspective results in a “recipe” for derving sequence layers by making three choices: the importance of each association $\{\gamma_{i}^{(t)}\}_{i=1}^{t}$, the regressor function class $\mathcal{M}$, and the optimization algorithm. Parametric regression layers tend to have an efficient recurrence for updating the memory ${m}_{t}$, while non-parametric layers like softmax-attention do not. For simplicity, we discuss causal sequence models where prefix key-value pairs are memorized, but the same principles also apply to non-causal ones (e.g. generically masked attention).
Section 2 Background on existing sequence modeling layers
The goal of sequence modeling is to transform a sequence of input tokens $\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$ into output tokens $\mathbf{y}_{1},\ldots,\mathbf{y}_{T}$ (Sutskever et al., 2014). A common design pattern, first proposed by Vaswani et al. (2017) is to transform the input tokens into key-value pairs $(\mathbf{k}_{1},\mathbf{v}_{1}),\ldots,(\mathbf{k}_{T},\mathbf{v}_{T})\in\mathbb{R}^{{D_{k}}}\times\mathbb{R}^{{D_{v}}}$ and queries $\mathbf{q}_{1},\ldots,\mathbf{q}_{T}\in\mathbb{R}^{D_{k}}$ before applying operations that “mix” the sequences together to produce the output tokens. Typically, each $\mathbf{q}_{t},\mathbf{k}_{t},$ and $\mathbf{v}_{t}$ is a linear transform of the corresponding input token $\mathbf{x}_{t}$, which can be thought of as different “views” of $\mathbf{x}_{t}$. Most subsequent sequence layers have kept this pattern but propose other ways of “mixing” queries, keys, and values. We begin by reviewing a few classes of existing sequence layers that follow this pattern before unifying them under our test-time regression framework.
Softmax attention.
Given input tokens $\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$, and their query, key, value vectors, softmax attention (Vaswani et al., 2017) computes each output token by
$\mathbf{y}_{t}=\sum_{i=1}^{t}\mathbf{v}_{i}\frac{s(\mathbf{k}_{i},\mathbf{q}_{t})}{\sum_{j=1}^{t}s(\mathbf{k}_{j},\mathbf{q}_{t})}$ (1)
for $t\in[T]$ where $s(\mathbf{k},\mathbf{q})=\exp(\mathbf{k}{}^{\top}\mathbf{q}/\sqrt{{D_{k}}})$. Each output $\mathbf{y}_{t}$ is a weighted sum of values $\{\mathbf{v}_{i}\}_{i=1}^{t}$, weighted by the similarity between the query $\mathbf{q}_{t}$ and the keys $\{\mathbf{k}_{i}\}_{i=1}^{t}$ associated with the values. This mechanism can be seen as a continuous-valued lookup table, using $\mathbf{q}_{t}$ to lookup an interpolated value for $\mathbf{y}_{t}$. Though highly effective, self-attention requires $\mathcal{O}\left(T^{2}\right)$ time to autoregressively compute the entire length $T$ sequence even after caching the keys and values, which can be prohibitively slow for generating long sequences.
Linear attention.
To address the computational challenges of self-attention, there has been a wealth of methods that seek to develop more efficient alternatives; see the survey by Tay et al. (2022). One particular class of more efficient sequence models has its origins in a recurrent variant of attention, known as linear attention (Katharopoulos et al., 2020). To achieve this efficiency, linear attention uses the standard inner product $\mathbf{k}{}^{\top}\mathbf{q}$, rather than using the exponentiated inner product $s(\mathbf{k}_{i},\mathbf{q}_{t})$, producing outputs of the form $\mathbf{y}_{t}\propto\sum_{i=1}^{t}\mathbf{v}_{i}\mathbf{k}_{i}{}^{\top}\mathbf{q}_{t}$ up to a normalization. This change allows all $T$ outputs to be recurrently computed in $\mathcal{O}\left(T\right)$ time, an asymptotic speedup, by updating a ${D_{v}}\times{D_{k}}$ “matrix-valued state” $\mathbf{M}_{t}=\sum_{i=1}^{t}\mathbf{v}_{i}\mathbf{k}_{i}{}^{\top}$, following the recurrence
$\mathbf{M}_{t}=\mathbf{M}_{t-1}+\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top},\quad\mathbf{y}_{t}=\mathbf{M}_{t}\mathbf{q}_{t}=\sum_{i=1}^{t}\mathbf{v}_{i}\mathbf{k}_{i}{}^{\top}\mathbf{q}_{t}.$ (2)
Each rank-1 outer product $\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top}$ can be thought of as writing a new key-value pair into memory. Although Katharopoulos et al. (2020)’s original formulation of linear attention included a normalization like softmax attention, we drop the normalization here since later works found that this normalization method causes training instabilities (Qin et al., 2022, Dao and Gu, 2024) .
To increase expressiveness, Katharopoulos et al. (2020) also proposed a variant of linear attention, inspired by the kernel trick (Schölkopf and Smola, 2001), that replaces the inner product $\mathbf{k}{}^{\top}\mathbf{q}$ with an inner product in a feature space $\phi(\mathbf{k}){}^{\top}\phi(\mathbf{q})$ where $\phi$ is a feature map $\phi:{D_{k}}\to{D_{\phi}}$. This modification results in the recurrence
$\mathbf{M}_{t}=\mathbf{M}_{t-1}+\mathbf{v}_{t}\phi(\mathbf{k}_{t}){}^{\top},\quad\mathbf{y}_{t}=\mathbf{M}_{t}\phi(\mathbf{q}_{t})=\sum_{i=1}^{t}\mathbf{v}_{i}\phi(\mathbf{k}_{i}){}^{\top}\phi(\mathbf{q}_{t}).$ (3)
Katharopoulos et al. (2020) used $\phi=1+\operatorname{ELU}$, but numerous alternative choices of feature maps have since been proposed, including the cosine function (Qin et al., 2021), deterministic parameter-free projections (Schlag et al., 2021), polynomial maps (Zhang et al., 2023a, Aksenov et al., 2024, Arora et al., 2024), random features (Peng et al., 2020, Choromanski et al., 2020, Chen et al., 2024), and more.
State-space models and linear attention with forgetting.
Since the recurrent state $\mathbf{M}_{t}$ of linear attention has finite size (${D_{v}}\times{D_{k}}$ numbers), $\mathbf{M}_{t}$ can only store a bounded amount of information. In Equation 2, $\mathbf{M}_{t}$ weighs all key value pairs equally across timesteps, trying to compress all key-value pairs equally well into its finite-sized state. As a result, the performance of linear attention can degrade as the sequence length increases. To address this limitation, various works have proposed adding a forgetting factor $\gamma_{t}$ to allow $\mathbf{M}_{t}$ to retain only the most recent information, decaying out key-value pairs over time. This produces the recurrence
$\mathbf{M}_{t}=\gamma_{t}\mathbf{M}_{t-1}+\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top},\quad\mathbf{y}_{t}=\mathbf{M}_{t}\mathbf{q}_{t}.$ (4)
At each step $t$, all previous values in $\mathbf{M}_{t-1}$ are attenuated by a factor of $\gamma_{t}\in[0,1]$ before the new value is stored. When $\gamma_{t}$ depends on $\mathbf{x}_{t}$, this forgetting factor can be seen as a forgetting gate, akin to that of an LSTM (Hochreiter and Schmidhuber, 1997). This class of linear attention with forgetting includes RetNet (Sun et al., 2023), gated linear attention (Yang et al., 2024b), Gateloop (Katsch, 2024), mLSTM (Beck et al., 2024), LRU (Orvieto et al., 2023), HGRN (Qin et al., 2024), and RWKV-6 (Peng et al., 2024). See Table 1 of Yang et al. (2024b) for how each model implements $\gamma_{t}$.
The forgetting factor in linear attention is also equivalent to another class of sequence layers known as state-space models (SSMs), which includes Mamba (Gu and Dao, 2024) and Mamba-2 (Dao and Gu, 2024). State-space models originated from a line of work on optimal online orthogonal polynomial projections (Gu et al., 2020, 2021, 2022) that enabled the architecture to retain information over long sequences. However, many recent works (Han et al., 2024, Trockman et al., 2024, Ren et al., 2024), including the Mamba-2 followup (Dao and Gu, 2024), have since shown that SSM layers are equivalent to linear attention with forgetting; the forgetting factor maps onto the time-varying step size and transition matrix used in SSMs.
Fast-weight layers and test-time training layers.
An alternative to attenuating all past information in $\mathbf{M}_{t-1}$ is to forget only the values whose keys are most similar to the new key $\mathbf{k}_{t}$. This was the motivation behind DeltaNet (Schlag et al., 2021), which is a revival of classic fast weight layers (Schmidhuber, 1992). DeltaNet follows a recurrence
$\mathbf{M}_{t}$ $=\mathbf{M}_{t-1}-\underbrace{(\mathbf{M}_{t-1}\mathbf{k}_{t})}_{\mathbf{v}_{t}^{\text{old}}}\mathbf{k}_{t}{}^{\top}+\underbrace{[\beta_{t}\mathbf{v}_{t}+(1-\beta_{t})\overbrace{\mathbf{M}_{t-1}\mathbf{k}_{t}}^{\mathbf{v}_{t}^{\text{old}}}]}_{\mathbf{v}_{t}^{\text{new}}}\mathbf{k}_{t}{}^{\top}$ (5)
$=\mathbf{M}_{t-1}(\mathbf{I}-\beta_{t}\mathbf{k}_{t}\mathbf{k}_{t}{}^{\top})+\beta_{t}\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top}$ (6)
where the first equation follows the notation of Schlag et al. (2021) in defining $\beta_{t}\in[0,1]$, $\mathbf{v}_{t}^{\text{old}}$, and $\mathbf{v}_{t}^{\text{new}}$. Like previous layers, the output at step $t$ is also $\mathbf{y}_{t}=\mathbf{M}_{t}\mathbf{q}_{t}.$ Comparing the recurrence of DeltaNet (Equation 5) to that of linear attention (Equation 2), we can interpret the update rule as rewriting the value that was bound to key $\mathbf{k}_{t}$. The update rule deletes an old value $\mathbf{v}_{t}^{\text{old}}=\mathbf{M}_{t-1}\mathbf{k}_{t}$ retrieved from the previous timestep’s memory $\mathbf{M}_{t-1}$ but using the new key $\mathbf{k}_{t}$. The layer then updates the recurrent state with a new interpolated value $\mathbf{v}_{t}^{\text{new}}$, binding it to $\mathbf{k}_{t}$.
The iterates $\mathbf{M}_{t}$ are also known as “fast-weights”, originally proposed by Schmidhuber (1992). These weights/iterates are “fast” because they change with each timestep, rapidly adapting to the input tokens, in contrast to the “slow-weights” in the rest of the neural network, such as the linear projections weights, which are only adjusted at training time. The traditional “slow-weights” can be seen as the weights of a metalearner (Clark et al., 2022), learning to use the “fast-weights” to learn from the input sequence (Sun et al., 2024).
Online learning layers.
The recurrences we have reviewed so far all share a common high-level behavior: at each timestep $t$, the recurrent state must be adapted to incorporate information from the latest key-value pair $(\mathbf{k}_{t},\mathbf{v}_{t})$. This property can be seen as making an online sequence of decisions, balancing between retaining historical information and adapting to new information.
Motivated by this sequential decision making perspective, Liu et al. (2024) proposed to derive recurrences by minimizing an online sequence of objectives:
$\mathbf{M}_{t}=\operatorname*{\arg\!\min}_{\mathbf{M}}\ d(\mathbf{M}_{t-1},\mathbf{M})+\ell_{t}(\mathbf{M})$ (7)
where $d(\mathbf{M}_{t-1},\mathbf{M})$ is a measure of the discrepancy between $\mathbf{M}$ and the previous iterate $\mathbf{M}_{t-1}$, encouraging retention of older information, and $\ell_{t}$ is a loss function that encourages adapting to new information. Once again, the output is computed by $\mathbf{y}_{t}=\mathbf{M}_{t}\mathbf{q}_{t}$. Recurrent layers derived via Equation 7 are called “online learning layers” or “online learners”. Recent examples of such layers include Longhorn (Liu et al., 2024), Gated DeltaNet(Yang et al., 2024a), and Titans (Behrouz et al., 2024), each of these layers was derived from specific choice of functions $d$ and $\ell_{t}$. See Table 2 from Yang et al. (2024a) for each layer’s choice of $d$ and $\ell_{t}$.
@@
Section 3 Test-time regression as a framework for designing sequence layers
Having discussed a number of existing sequence layers, each motivated by an atomic architectural modification, we now present a unified framework to clarify their relationship to each other. Our approach is motivated by empirical findings that a model’s associative recall ability strongly correlates with its performance and in-context learning capabilities (Olsson et al., 2022, Arora et al., 2023a). We formalize the two step process of memorization and retrieval, presenting a principled recipe for deriving sequence layers with mathematically grounded recall abilities.
Step 1: Memorization as regression.
@@
From our everyday experience, we already have an intuition for how associative memory should behave; for example, hearing a friend’s name should trigger a mental impression of that friend. This pairing of a “cue” and a “response” is known as an association. To reflect terminology in deep learning (Vaswani et al., 2017), we will refer to cues as “keys” and responses as “values”. With this intuition, we can define a mathematical model for associative memory. Given key-value pairs $(\mathbf{k}_{1},\mathbf{v}_{1}),\ldots,(\mathbf{k}_{t},\mathbf{v}_{t})\in\mathbb{R}^{D_{k}}\times\mathbb{R}^{D_{v}}$ to be memorized, an associative memory system is a function ${m}:\mathbb{R}^{D_{k}}\to\mathbb{R}^{D_{v}}$ such that ${m}(\mathbf{k}_{i})\approx\mathbf{v}_{i}$ for $i=1,\ldots,t$ (Kohonen, 1972, 1989). Such a mapping ${m}$ between two sets of related objects is also known as a hetero-associative memory in classic signal processing and neurocomputing literature (Hinton and Anderson, 1989).
@@
Our main observation is that finding a memory map that satisfies these constraints reduces to finding a vector-valued regressor ${m}$. We can then store a set of key-value associations into an associative memory system by solving the weighted regression problem
[…]
@@
where the importance of each association is controlled by adjusting the relative weights $\{\gamma_{i}^{(t)}\}_{i=1}^{t}$ Classic works on associative memory are restricted to linear memory maps, corresponding to linear regression (Kohonen, 1989, Chapter 6.6), but here we allow $\mathcal{M}$ to be any function class. We refer to this relationship between regression and associative memory as the “regression-memory correspondence”.
Step 2: Memory retrieval as function application.
Having derived a regression-based procedure to memorize a set of associations into ${m}$, we can now retrieve values from memory by applying ${m}$ to a new vector, which we call a “query” $\mathbf{q}$, producing output
[…]
The retrieval performance depends on how well Equation 8 is minimized. When Equation 8 attains zero-loss, ${m}_{t}$ memorizes all associations perfectly, and we get perfect retrieval: ${m}_{t}(\mathbf{k}_{i})=\mathbf{v}_{i}$ for $i=1,\ldots,t$. For certain function classes $\mathcal{M}$, e.g. smooth functions, ${m}$ can also extrapolate outside of the existing set of keys. In those cases, ${m}_{t}(\mathbf{q})$ instead retrieves a value based on the proximity of $\mathbf{q}$ to the existing set of keys.
Sequence layers that memorize then retrieve.
Having formalized memorization and retrieval via our regression framework, we can now combine the two to produce sequence layers that directly performs associative recall as a forward pass. Recall that we are given $\{\mathbf{q}_{i}\}_{i=1}^{T}$, $\{\mathbf{k}_{i}\}_{i=1}^{T}$, and $\{\mathbf{v}_{i}\}_{i=1}^{T}$, and we seek to produce a set of outputs $\{\mathbf{y}_{i}\}_{i=1}^{T}$ with our sequence layer. The forward pass of a test-time regression layer first memorizes a set of key-value pairs and then produces $\mathbf{y}_{t}$ by retrieving a value from its associative memory. More specifically, in causal modeling, at each timestep $t$, we first solve Equation 8 to get a regressor ${m}_{t}$ that memorizes the prefix $\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i=1}^{t}$; then we apply ${m}_{t}$ to $\mathbf{q}_{t}$ to get $\mathbf{y}_{t}={m}_{t}(\mathbf{q}_{t})$. This naturally generalizes to non-causal tasks where ${m}_{t}$ can depend on suffix key-value pairs $\{\mathbf{k}_{i},\mathbf{v}_{i}\}_{i>t}$ as well. Our procedure is compatible with multihead design patterns inspired by multihead attention (Vaswani et al., 2017), multi-query attention (Shazeer, 2019), and grouped query attention (Ainslie et al., 2023). For example, multihead test-time regression can be implemented using $H$ independent sets of queries, keys, and values, resulting in $H$ independent associative memory maps at each timestep.
A recipe for designing your own sequence layer.
@@
Although Equation 8 succinctly summarizes the memorization step of our framework, it leaves many details unspecified. For example, over which class of regressors $\mathcal{M}$ should we consider? Additionally, which associations should we prioritize memorizing? Finally, which optimization procedure should we use to minimize the loss? The regression-memory correspondence thus allows for a vast design space. We partition this wide landscape into three simple design choices:
@@
To demonstrate how our framework unifies our understanding of sequence models, we succinctly derive a swath of recently proposed sequence layers. We categorize existing sequence layers into four distinct classes: linear attention (possibly with feature maps), linear attention with gating (which includes state-space models like Mamba), fast-weight layers, online learners, and softmax self-attention.